# Scoring Nim

### v1

@@ Abstract
Nim is a well-known combinatorial game, in which two players alternately remove stones from distinct piles. A player who removes the last stone wins under the normal play rule, while a player loses under the misère play rule. In this paper, we propose a new variant of Nim with scoring that generalizes both the normal and misère play versions of Nim as special cases. We study the theoretical aspects of this extended game and analyze its fundamental properties, such as optimal strategies and payoff functions.
@@ 1 Introduction
Nim is a well-known combinatorial game that has been extensively studied. In the standard version of Nim, there are three piles of stones, and two players take turns. On each turn, a player selects one pile and removes any positive number of stones from that pile. The position of the game is represented as $(x,y,z)$ where $x,y$ and $z$ are non-negative integers representing the number of stones in each pile. For example, from $(3,5,7)$, the player whose turn it is can move to $(3,2,7)$, $(0,5,7)$, and so on. Players take turns making these moves, and whoever takes the last stone wins the game, while a player with no available moves automatically loses. This rule is known as the normal play rule. If the position is $(3,0,0)$, the player whose turn it is can move to $(0,0,0)$, the terminal position, and win the game. This game can be played with any number of piles following the same rules.
In contrast, The rule where the player who takes the last stone loses is known as the misère play rule. Under this rule, if the position is $(3,0,0)$, the player whose turn it is can move to $(1,0,0)$, thereby forcing the opponent to take the last stone and securing victory.
Nim is a zero-sum, deterministic, perfect-information, impartial and finite game with no possibility of a draw. Therefore, every position can be classified as either an N-position, where the player whose turn it is can force a win, or a P-position, where the player whose turn it is will inevitably lose if the opponent plays optimally. The classification of N-positions and P-positions in Nim was first established by Bouton [3] in 1902, revealing its mathematical structure. Nim and its variants are discussed in detail for instance, by Albert, Nowakowski and Wolfe [1], Berlekamp, Conway and Guy [2] and Siegel [6].
@@
This paper is organized as follows. In Section 2 we formally introduces the rules of the new game. In Section 3 we define the payoff function, which plays a crucial role in the analysis of this game. In Section 4 we examine the optimal strategy for the case of two piles of stones. In Section 5 we discuss general properties of the payoff function. In Section 6 we fix the number of piles at three and explicitly present the payoff function when the minimum number of stones in a pile is one. We analyze the reasons behind the numerous breakpoints (i.e., points where the increasing and decreasing trends switch) in the plot of the function.
@@ 2 The Rules of Scoring Nim
  • $1$ point for each stone taken,
  • $N$ points for taking the last stone.
@@
A game position is represented [^1] When the number of piles is 1, 2, or 3, the game positions are represented as $(x)$, $(x,y)$, and $(x,y,z)$, respectively. as ${\bm{p}}=(p_{1},p_{2},\ldots,p_{n})$, where each $p_{i}$ is a non-negative integer. The terminal position $(0,0,\ldots,0)$ is denoted by $\bm{0}$. The total number of stones in a position $\bm{p}$ is denoted as $|\bm{p}|=\sum_{i=1}^{n}p_{i}$. As a concrete example, consider a game where the initial position is $(3,2,1)$, and two players (the first and second player) take turns as follows:
@@
Furthermore, Scoring Nim becomes even more intriguing when the objective shifts from merely securing victory to maximizing points. This setting corresponds to scoring games in combinatorial game theory (for details, see Stewart [7].). In [7], Section 5 primarily discusses the scoring adaptation of octal games as a topic related to Nim [^2] Additionally, in the introduction of [7], there is a brief mention of a stone-taking game corresponding to Scoring Nim with $N=0$.. Scoring Nim provides a simple yet fascinating extension of Nim by introducing scores. It serves as a valuable example of scoring games and offers important insights into the study of such games.
Let us consider Scoring Nim with an initial position of $(5,4,2)$. Players’ move selections at each turn can be represented as an extensive-form game tree. By applying the minimax method, one can determine the optimal move for each position and the optimal sequence of moves leading to the terminal position, assuming both players play optimally. Figure 1 presents the game tree for the initial sequence of moves, illustrating some representative promising moves.
$(5,4,2)$$(5,4,1)$$(1,4,2)$$(0,4,2)$$(2,4,1)$$(1,3,2)$$(1,0,2)$$(0,2,2)$$(0,0,2)$
Figure 1: A partial game tree from the initial position $(5,4,1)$
In this initial position, the optimal move for the first player is:
@@
  • Moves to $(1,4,2)$ or $(0,4,2)$ when $-4\leq N\leq-1$,
  • Moves to $(0,4,2)$ when $-1\leq N\leq 1$,
@@
In this case, the first player takes $8$ stones and earns $8$ points, while the second player takes $3$ stones and, with the bonus, obtains $63+N=$ 6 points. Intuitively, these intermediate strategies can be understood as moves that force the opponent into a position where neither the Nim-winning strategy nor the greedy strategy is particularly effective. A detailed discussion of these intermediate strategies for such values of $N$ will be provided in Section 6.
@@ 3 The Payoff Function and Recursion Formula
Scoring Nim is a constant-sum game in which the total points of both players always add up to $|\bm{p}|+N$ when starting from the position $\bm{p}$. Therefore, treating it as a zero-sum game, we assume that each player aims to maximize the final score difference, i.e.,
[…]
@@
at the end of the game. The sign of this value indicates whether the player has gained more points than the opponent. Let $f_{N}(p_{1},p_{2},\ldots,p_{n})$ denote this score difference for the first player, given in the initial position $\bm{p}=(p_{1},p_{2},\ldots,p_{n})$. Then, for $n=1$ (i.e., when there is only one pile of stones), we obtain
$f_{N}(1)$ $=$ $1+N,$
$f_{N}(x)$ $=$ $\max\{x+N,x-2-N\}=x-1+|1+N|$ (1)
@@
for $x=1$ and $x\geq 2$, respectively. From the comparison in (1), when the initial position is $(x)$ with $x\geq 2$, the first player should take all the stones if $N\geq-1$. On the other hand, if $N\leq-1$, the first player should take $x-1$ stones on the first move, leaving only one stone for the opponent’s turn.
@@
Similar to $f_{N}(x)$, $f_{N}(\bm{p})$ is recursively defined based on comparisons with subsequent positions. To formalize this, we extend the definition of the payoff function $f_{N}(\bm{p})$ to positions beyond the initial state. we define it as the difference in total points accumulated by the two players from the position forward, for the current player. To ensure consistency, we set $f_{N}(\bm{0})=-N$, meaning that when no stones remain, the opponent has just received an endgame bonus of $N$ points. Let $\bm{p}\to\bm{p}^{\prime}$ denote that a transition from position $\bm{p}$ to $\bm{p}^{\prime}$ is possible in a single move. Then, for a position $\bm{p}$ that is not a terminal position, the following recurrence relation holds:
$f_{N}(\bm{p})=\max_{\bm{p}\to\bm{p}^{\prime}}\{|\bm{p}|-|\bm{p}^{\prime}|-f_{N}(\bm{p}^{\prime})\}.$ (2)
@@
Here, $|\bm{p}|-|\bm{p}^{\prime}|$ means the number of stones taken in the move from $\bm{p}$ to $\bm{p}^{\prime}$, with each stone contributing one point. This recursion formula allows us to determine $f_{N}(\bm{p})$ sequentially, starting from positions with a smaller total number of stones.
The payoff function $f_{N}(\bm{p})$ is a function of the position $\bm{p}$, but if we fix $\bm{p}$ and regard the bonus point $N$ as a variable, it can also be considered a function of $N$ with the real numbers as its domain. In this case, from the recursion formula (2) and the fact that $f_{N}(\bm{0})$ is a continuous function with a slope of $-1$, the following properties can be derived.
Proposition 3.1.
When considering $f_{N}(\bm{p})$ as a function of $N$, the following hold:
(i) $f_{N}(\bm{p})$ is continuous.
(ii) If $N$ is an integer, then $f_{N}(\bm{p})$ is also an integer, and its parity coincides with that of $|\bm{p}|+N$.
(iii) When $N$ is not an integer, the slope of $f_{N}(\bm{p})$ is either $1$ or $-1$. Furthermore, if the slope is $1$, the first player takes the last stone under both players’ optimal strategies, whereas if the slope is $-1$, the second player takes the last stone under both players’ optimal strategies.
Since this proposition is trivial, the proof is omitted. By this proposition, when considering $f_{N}(\bm{p})$ as a function of $N$, if the values for integer $N$ are determined, then the values for all real $N$ can also be obtained by linear interpolation.
The following fact, which follows from Proposition 3.1(iii), is also useful and is presented here as a corollary.
Corollary 3.2.
For real numbers $N$ and $N^{\prime}$,
$|f_{N}(\bm{p})-f_{N^{\prime}}(\bm{p})|\leq|N-N^{\prime}|$
holds.
@@
As a concrete example of this payoff function, we illustrate $f_{N}(5,4,2)$ as a function of $N$ in Figure 2 for the game that was used in the previous section. As an example of a similar position, we will present the plot of the payoff function $f_{N}(5,4,3)$ later in Remark 6.5 and Figure 6.
[…]
@@ 4 Scoring Nim with Two Piles
In this section, we derive an explicit formula for $f_{N}(\bm{p})$ when the initial position has two piles. Using equations (1) and (2), we first obtain the following results.
Proposition 4.1.
For $x\geq 2$,
$f_{N}(1,1)=-N,\quad f_{N}(x,1)=x-1+|N|.$
Proof. In the position $(1,1)$, each player takes one stone; the first player earns 1 point, while the second player earns $1+N$ points. Moreover, applying the recursion formula (2),
$f_{N}(2,1)=\max\{2-(1+N),1+N,1-(1+|1+N|)\}=1+|N|$
holds, and for $x\geq 3$, we derive $f_{N}(x,1)$ similarly.  $\Box$
In the position $(x,1)$ with $x\geq 2$, the optimal strategy depends on the value of $N$. If $N\geq 0$, the optimal move is to take $x-1$ stones, leading to the position $(1,1)$. On the other hand, if $N\leq 0$, optimal move is to take $x$ stones, leading to the position $(0,1)$. Since the number of possible moves is relatively small for the two-pile case, $f_{N}(x,y)$ can be computed directly using the recursion formula. The following proposition completes the analysis for all cases.
Proposition 4.2.
The function $f_{N}(x,y)$ is defined as follows when $x,y\geq 2$.
$f_{N}(x,x)=x-f_{N}(0,x)=1-|1+N|.$ (3)
(ii) When $x>y$,
$f_{N}(x,y)$ $=$ $\max\{x-f_{N}(0,y),x-y-f_{N}(y,y)\}$ (4)
$=$ $x-y\ +\bigm{|}1-|1+N|\bigm{|}.$
If both piles contain the same number of stones, the optimal move is to take all stones from one pile. If the number of stones in the two piles is different, if $-2\leq N\leq 0$, the optimal move follows the greedy strategy, which consists of taking all stones from the larger pile. Otherwise, the optimal move follows the Nim-winning strategy by equalizing the number of stones in both piles. The proof follows the same approach as in Proposition 4.1 and can be established by induction on the total number of stones, using the recursion formula. These results indicate that Scoring Nim with two piles has well-defined strategic properties.

### v2

@@ Abstract
Nim is a well-known combinatorial game in which two players alternately remove stones from distinct piles. A player who removes the last stone wins under the normal play rule, while a player loses under the misère play rule. In this paper, we propose a new variant of Nim with scoring that generalizes both the normal and misère play versions of Nim as special cases. We study the theoretical aspects of this extended game and analyze its fundamental properties, such as optimal strategies and payoff functions.
@@ 1 Introduction
Nim is a well-known combinatorial game that has been extensively studied. In the standard version of Nim, there are three piles of stones, and two players take turns. On each turn, a player selects one pile and removes any positive number of stones from that pile. The position of the game is represented as $(x,y,z)$ where $x,y$ and $z$ are non-negative integers representing the number of stones in each pile. For example, from $(3,5,7)$, the player to move can move to $(3,2,7)$, $(0,5,7)$, and so on. Players take turns making these moves, and whoever takes the last stone wins the game, while a player with no available moves automatically loses. This rule is known as the normal play rule. If the position is $(3,0,0)$, the player to move can move to $(0,0,0)$, the terminal position, and win the game. This game can be played with any number of piles following the same rules.
In contrast, the rule where the player who takes the last stone loses is known as the misère play rule. Under this rule, if the position is $(3,0,0)$, the player to move can move to $(1,0,0)$, thereby forcing the opponent to take the last stone and securing victory.
Nim is a zero-sum, deterministic, perfect-information, impartial and finite game with no possibility of a draw. Therefore, every position can be classified as either an N-position, where the player to move can force a win, or a P-position, where the player to move will inevitably lose if the opponent plays optimally. The classification of N-positions and P-positions in Nim was first established by Bouton (1901/1902) Bouton (1901/02), revealing its mathematical structure. Nim and its variants are discussed in detail for instance by Albert, Nowakowski and Wolfe (2011) Albert and Wolfe (2011), Berlekamp, Conway and Guy (2001, 2003, 2004) Berlekamp et al. (2001,2003,2004) and Siegel (2013) Siegel (2013).
@@
This paper is organized as follows. In Section 2 we formally introduce the rules of the new game. In Section 3 we define the payoff function, which plays a crucial role in the analysis of this game. In Section 4 we examine the optimal strategy for the case of two piles of stones. In Section 5 we discuss general properties of the payoff function. In Section 6 we fix the number of piles at three and explicitly present the payoff function when the minimum number of stones in a pile is one. We analyze the reasons behind the numerous breakpoints (i.e., points where the increasing and decreasing trends switch) in the plot of the function.
This paper is a compilation of Oginuma and Shinoda (2024) Oginuma and Shinoda (2024a, b). This resubmission is allowed under the guidelines of the Information Processing Society of Japan (For details, see
https://www.ipsj.or.jp/english/faq/ronbun-faq_e.html).
@@ 2 The rules of Scoring Nim
  • 11 point for each stone taken,
  • NN points for taking the last stone.
@@
A game position is represented [^1] When the number of piles is 1, 2, or 3, the game positions are represented as $(x)$, $(x,y)$, and $(x,y,z)$, respectively. as ${\boldsymbol{p}}=(p_{1},p_{2},\ldots,p_{n})$, where each $p_{i}$ is a non-negative integer. The terminal position $(0,0,\ldots,0)$ is denoted by $\boldsymbol{0}$. The total number of stones in a position $\boldsymbol{p}$ is denoted as $|\boldsymbol{p}|=\sum_{i=1}^{n}p_{i}$. As a concrete example, consider a game where the initial position is $(3,2,1)$, and two players (the first and second player) take turns as follows:
@@
Furthermore, Scoring Nim becomes even more intriguing when the objective shifts from merely securing victory to maximizing points. This setting corresponds to scoring games in combinatorial game theory (for details, see Stewart (2019) Stewart (2019)). In Stewart (2019), Section 5 primarily discusses the scoring adaptation of octal games as a topic related to Nim [^2] Additionally, in the introduction of Stewart (2019), there is a brief mention of a stone-taking game corresponding to Scoring Nim with $N=0$.. Scoring Nim provides a simple yet fascinating extension of Nim by introducing scores. It serves as a valuable example of scoring games and offers important insights into the study of such games. While our analysis does not explicitly employ the general theoretical framework of scoring play, the game we introduce falls naturally within the class of impartial scoring games considered in the literature, and may provide a useful concrete case for further theoretical investigation.
Let us consider Scoring Nim with an initial position of $(5,4,2)$. Players’ move selections at each turn can be represented as an extensive-form game tree. By applying the minimax method, one can determine the optimal move for each position and the optimal sequence of moves leading to the terminal position, assuming both players play optimally. Fig. 1 presents the game tree for the initial sequence of moves, illustrating some representative promising moves.
(5,4,2)(5,4,2)(5,4,1)(5,4,1)(1,4,2)(1,4,2)(0,4,2)(0,4,2)(2,4,1)(2,4,1)(1,3,2)(1,3,2)(1,0,2)(1,0,2)(0,2,2)(0,2,2)$(0,0,2)$
Figure 1: A partial game tree from the initial position $(5,4,2)$
The following conclusions are presented here as preliminary observations in order to illustrate the dependence of optimal play on the value of the bonus $N$; a general theoretical framework justifying them will be established in subsequent sections. In this initial position, the optimal move for the first player is:
@@
  • Moves to $(1,4,2)$ or $(0,4,2)$ when $4\leq N\leq-1$,
  • Moves to $(0,4,2)$ when $1\leq N\leq 1$,
@@
In this case, the first player takes $8$ stones and earns $8$ points, while the second player takes $3$ stones and, with the bonus, obtains $3+N=6$ points. Intuitively, these intermediate strategies can be understood as moves that force the opponent into a position where neither the Nim-winning strategy nor the greedy strategy is particularly effective. A detailed discussion of these intermediate strategies for such values of $N$ will be provided in Section 6.
@@ 3 The payoff function and recursion formula
Scoring Nim is a constant-sum game in which the total points of both players always add up to $|\boldsymbol{p}|+N$ when starting from the position $\boldsymbol{p}$. Therefore, treating it as a zero-sum game, we assume that each player aims to maximize the final score difference, i.e.,
$\mbox{Payoff}=(\mbox{First player's score})-(\mbox{Second player's score})$
@@
at the end of the game. The sign of this value indicates whether the player has gained more points than the opponent. Let $f_{N}(p_{1},p_{2},\ldots,p_{n})$ denote this score difference for the first player, given in the initial position $\boldsymbol{p}=(p_{1},p_{2},\ldots,p_{n})$. Then, for $n=1$ (i.e., when there is only one pile of stones), we obtain $f_{N}(1)=1+N$ and
$f_{N}(x)=\max\{x+N,x-2-N\}=x-1+|1+N|$ (3.1)
@@
for $x\geq 2$. From the comparison in (3.1), when the initial position is $(x)$ with $x\geq 2$, the first player should take all the stones if $N\geq-1$. On the other hand, if $N\leq-1$, the first player should take x1x-1 stones on the first move, leaving only one stone for the opponent’s turn.
@@
Similar to $f_{N}(x)$ with $x\geq 2$, $f_{N}(\boldsymbol{p})$ is recursively defined based on comparisons with subsequent positions. To formalize this, we extend the definition of the payoff function $f_{N}(\boldsymbol{p})$ to positions beyond the initial state. We define it as the difference between the total points obtained by the current player and those obtained by the opponent from that position onward. To ensure consistency, we set $f_{N}(\boldsymbol{0})=-N$, meaning that when no stones remain, the opponent has just received an endgame bonus of $N$ points. Let $\boldsymbol{p}\to\boldsymbol{p}^{\prime}$ denote that a transition from position $\boldsymbol{p}$ to $\boldsymbol{p}^{\prime}$ is possible in a single move. Then, for a position $\boldsymbol{p}$ that is not a terminal position, the following recurrence relation holds:
$f_{N}(\boldsymbol{p})=\max_{\boldsymbol{p}\to\boldsymbol{p}^{\prime}}\{|\boldsymbol{p}|-|\boldsymbol{p}^{\prime}|-f_{N}(\boldsymbol{p}^{\prime})\}.$ (3.2)
@@
Here, $|\boldsymbol{p}|-|\boldsymbol{p}^{\prime}|$ means the number of stones taken in the move from $\boldsymbol{p}$ to $\boldsymbol{p}^{\prime}$, with each stone contributing one point. This recursion formula allows us to determine $f_{N}(\boldsymbol{p})$ sequentially, starting from positions with a smaller total number of stones.
@@
The payoff function $f_{N}(\boldsymbol{p})$ is a function of the position $\boldsymbol{p}$, but if we fix $\boldsymbol{p}$ and regard the bonus point $N$ as a variable, it can also be considered a function of $N$ with the real numbers as its domain. To illustrate the payoff function, Fig. 2 shows $f_{N}(5,4,2)$ as a function of $N$ for the game that was used in the previous section. A similar position, $f_{N}(5,4,3)$, will be presented later in Remark 1 and Fig. 6 in Section 6.
0334- [^4] NN
@@
The payoff function $f_{N}(\boldsymbol{p})$ satisfies the following properties.
Proposition 1.
(i) If $N$ is an integer, then $f_{N}(\boldsymbol{p})$ is also an integer, and its parity coincides with that of $|\boldsymbol{p}|+N$.
(ii) When $f_{N}(\boldsymbol{p})$ is regarded as a function of $N$, it is continuous in $N$, and for non-integer values of $N$, its slope is either $1$ or 1-1. In the former case, the first player takes the last stone under optimal play, whereas in the latter case, the second player does.
Proof. (i) If $N$ is an integer, then the scores of the first and second players are integers, and their sum is $|\boldsymbol{p}|+N$. Observe that
$f_{N}(\boldsymbol{p})=|\boldsymbol{p}|+N-2(\mbox{Second player's score}),$
and hence the parity coincides with that of $|\boldsymbol{p}|+N$.
(ii) We prove the claims by induction on the total number of stones $|\boldsymbol{p}|$. First consider $|\boldsymbol{p}|=0$, that is, the terminal position. In this case $f_{N}(\boldsymbol{0})=-N$, which is a continuous function of $N$ with slope 1-1. Since this position can be regarded as the one obtained when the previous player takes the last stone and the game ends, assertion (ii) holds.
Assume that the statements hold for all positions with $|\boldsymbol{p}|<m$. For $|\boldsymbol{p}|=m$, the recursion formula (3.2) shows that $f_{N}(\boldsymbol{p})$ is defined as the maximum of finitely many continuous functions, each having slope $\pm 1$. Hence it is continuous and its slope is either $1$ or 1-1. Moreover, for a non-integer $N$, whenever there exists a position $\boldsymbol{p^{\prime}}$ such that $f_{N}(\boldsymbol{p})=|\boldsymbol{p}|-|\boldsymbol{p^{\prime}}|-f_{N}(\boldsymbol{p^{\prime}})$, the slope of $f_{N}(\boldsymbol{p})$ has the opposite sign to that of $f_{N}(\boldsymbol{p^{\prime}})$. This corresponds to the alternation of turns, and therefore the player who takes the last stone is determined according to this sign.  $\Box$
By this proposition, when considering $f_{N}(\boldsymbol{p})$ as a function of $N$, if the values for integer $N$ are determined, then the values for all real $N$ can also be obtained by linear interpolation.
The following fact follows from Proposition 1 (ii), since it implies that the absolute value of the slope of $f_{N}(\boldsymbol{p})$ is at most 1. As this fact is useful, we present it here as a corollary.
Corollary 2.
For real numbers $N$ and $N^{\prime}$, $|f_{N}(\boldsymbol{p})-f_{N^{\prime}}(\boldsymbol{p})|\leq|N-N^{\prime}|$ holds.
@@ 4 Scoring Nim with two piles
In this section, we derive an explicit formula for $f_{N}(\boldsymbol{p})$ when the initial position has two piles. Using equations (3.1) and (3.2), we first obtain the following results.
Proposition 3.
$f_{N}(1,1)=-N$, and
$\quad f_{N}(x,1)=x-1+|N|$
for $x\geq 2$.
Proof. In the position $(1,1)$, each player takes one stone; the first player earns 1 point, while the second player earns $1+N$ points. Moreover, applying (3.1) and the recursion formula (3.2),
$f_{N}(x,1)$
$=$ $\max\left\{\begin{array}[]{l}x-f_{N}(0,1),x-1-f_{N}(1,1),x-2-f_{N}(2,1),\ldots,\\ 1-f_{N}(x-1,1),1-f_{N}(x,0)\end{array}\right\}$
$=$ $\max\left\{\begin{array}[]{l}x-1-N,x-1+N,x-3-|N|,\ldots,\\ -x+3+|N|,-x+2-|1+N|\end{array}\right\}$
$=$ $x-1+|N|$
holds for $x\geq 2$ inductively.  $\Box$
In the position $(x,1)$ with $x\geq 2$, the optimal strategy depends on the value of $N$. If $N\geq 0$, the optimal move is to take x1x-1 stones, leading to the position $(1,1)$. On the other hand, if $N\leq 0$, the optimal move is to take $x$ stones, leading to the position $(0,1)$. Since the number of possible moves is relatively small for the two-pile case, $f_{N}(x,y)$ can be computed directly using the recursion formula. The following proposition completes the analysis for all cases.
Proposition 4.
Let $x,y\geq 2$. Then $f_{N}(x,y)$ satisfies the following.
$f_{N}(x,x)=x-f_{N}(0,x)=1-|1+N|.$ (4.3)
(ii) If $x>y$,
$f_{N}(x,y)=\max\{x-f_{N}(0,y),x-y-f_{N}(y,y)\}=x-y+|1-|1+N||.$ (4.4)
If both piles contain the same number of stones, the optimal move is to take all stones from one pile. When the piles have different sizes and $2\leq N\leq 0$, the optimal move follows the greedy strategy of taking all stones from the larger pile. Otherwise, the optimal move follows the Nim-winning strategy of equalizing the numbers of stones in the two piles. The proof follows the same approach as in Proposition 3 and can be established by induction on the total number of stones, using the recursion formula.
Proof. We prove (i) and (ii) for all $(x,y)$ satisfying $x\geq y\geq 2$ by induction. Assume that (4.3) and (4.4) hold for every $(x^{\prime},y^{\prime})$ with $2\leq x^{\prime}\leq x$, $2\leq y^{\prime}\leq y$ and $x^{\prime}+y^{\prime}<x+y$. Then, together with (3.1) and Proposition 3, it follows that the following equalities hold:
$f_{N}(x,x)$
$=$ $\max\{x-f_{N}(x,0),x-1-f_{N}(x,1),x-2-f_{N}(x,2),\ldots,1-f_{N}(x,x-1)\}$
$=$ $\max\{1-|1+N|,-|N|,-|1-|1+N||,\ldots,-|1-|1+N||\}$
$=$ $1-|1+N|,$
$f_{N}(x,y)$
$=$ $\max\left\{\begin{array}[]{l}x-f_{N}(0,y),x-1-f_{N}(1,y),\ldots,x-y-f_{N}(y,y),\ldots,\\ 1-f_{N}(x-1,y),\\ y-f_{N}(x,0),y-1-f_{N}(x,1),y-2-f_{N}(x,2),\ldots,\\ 1-f_{N}(x,y-1)\end{array}\right\}$
$=$ $\max\left\{\begin{array}[]{l}x-y+1-|1+N|,x-y-|N|,\ldots,x-y-1+|1+N|,\ldots,\\ -x+y+2-|1-|1+N||,\\ -x+y+1-|1+N|,-x+y-|N|,-x+y+|1-|1+N||,\ldots,\\ -x+y+|1-|1+N||\end{array}\right\}$
$=$ $x-y+|1-|1+N||.$
In the last maximization step, we use the assumption that $x>y$.  $\Box$
These results indicate that Scoring Nim with two piles has well-defined strategic properties.